If you add 600 cal of heat to 50 g of water at 10oC, what will be the temperature of the water in oC?

1
Expert's answer
2021-11-23T10:31:24-0500

"600\ cal=2510.4\ J"


"Q=cm(t-t_0)\to t=Q/(cm)+t_0=2510.4/(4182\cdot0.05)+10=22°C" . Answer